The planet Mars has a mass of 6.39 x 10^23kg and a radius of 3.39 x 10^6m the gravitational field strength on the surface of Mars is______N/kg?
1
Expert's answer
2020-06-08T10:12:37-0400

We can find the gravitational field strength  on the surface of Mars from the formula:


"g_{Mars} = \dfrac{GM_{Mars}}{r_{Mars}^2},""g_{Mars} = \dfrac{6.67 \cdot 10^{-11} \ \dfrac{N \cdot m^2}{kg^2} \cdot 6.39 \cdot 10^{23} \ kg}{(3.39 \cdot 10^6 \ m)^2} = 3.71 \ \dfrac{N}{kg}."

Answer:

"g_{Mars} = 3.71 \ \dfrac{N}{kg}."